Someone bought a CD player for $250 and finds 2 years later when he wants to sell it that it's worth $100. Find a formula for the value of the CD player if the value decreases.
Answer by stanbon(75887)

Someone bought a CD player for $250 and finds 2 years later when he wants to sell it that it's worth $100. Find a formula for the value of the CD player if the value decreases.
============
You have 2 points: (0,250) and (2,100)
slope = (100-250)/(2-0) = -150/2 = -75
------
Form: y = mz + b
250 = -75*0 + b
b = 250
------
Equation:
value = -75(years) + 250
================================
Cheers,
Stan H.
=================